# Heterogeneous User Modeling for LLM-based Recommendation

### v1

@@ 3.2. Robustness Enhancer
Domain smoothing. Directly applying domain importance during training may cause instability, particularly for sparse domains with few samples (Wen et al., 2022). Due to the scarcity of training samples, the estimated losses for these domains can exhibit high variance, leading to noisy or unreliable values that do not accurately reflect their true importance. To alleviate this, we adopt a domain smoothing strategy following (Wen et al., 2022; Lin et al., 2024e), which incorporates prior information into the current domain importance estimation. We maintain a domain importance vector and introduce a regulating factor to control the update smoothness. This leads to the following optimization objective:
@@ 4.3.1. Ablation Study (RQ2)
[…]
@@
4.3.2. Robustness on seen domains (RQ2)
@@ 4.3.3. Noise resistance(RQ2)
4.3.4. Generalization on unseen domains (RQ3)
@@ 4.3.6. Effect of mask ratio
To investigate the impact of the mask ratio on the accuracy of HUM, we vary the ratio from to and present the results in Figure 9. It is observed that 1) Increasing within a small range can improve model performance. This is because the masking mechanism allows the model to acquire useful heterogeneous information from other domains, thereby enhancing recommendation performance in the target domain. Additionally, it potentially filters out the impact of noisy interactions, further improving recommendation accuracy. 2) However, excessively increasing may inversely hurt performance. This could be because an overly high mask ratio leads to information loss, resulting in suboptimal learning of target domain information. 3) These observations suggest that there is a sweet spot for , where the benefits of cross-domain information outweigh the costs of missing target domain details.
@@ 5. Related Work
LLM-based Recommendation. Large Language Models (LLMs) have shown exceptional success in various recommendation tasks (Wang et al., 2024; Fu et al., 2024; Ma et al., 2025; Bao et al., 2023; Deng et al., 2025). There are two primary methods for using LLMs in recommendation: 1) LLM-based recommender, that directly employs LLMs as the recommendation models (Li et al., 2024; Zhang et al., 2023; Lin et al., 2024a; Bao et al., 2023; Lin et al., 2024b; Prakash et al., 2023; Zheng et al., 2024). 2) LLM-enhanced recommender, while utilize LLMs for data augmentation (Wei et al., 2024; Xi et al., 2023; Liu et al., 2024) and representation (Qiu et al., 2021; Ren et al., 2024; Li et al., 2023; Tang et al., 2023). Leveraging LLMs for representation refers to generating embedding for entities(e.g., users or items) through LLMs. For instance, LLM-Rec (Tang et al., 2023) uses BERT (Devlin et al., 2018), OPT (Zhang et al., 2022) and Flan-T5 (Chung et al., 2024) to create user and item embedding based on their semantic information(i.e., plain text). In this work, we focus on leveraging LLMs for representing heterogeneous user interactions, thus we do not dive deeper into the discussion of representation methods in other fields of recommendation.
@@ 6. Conclusion and Future Work
This work emphasizes the importance of heterogeneous user modeling in LLM-based open-domain recommendation and suggests several promising directions for future research: 1) Future work could focus on integrate temporal dynamics into the user modeling process, considering how user preferences and behaviors evolve over time. 2) Another direction is to explore how to combine heterogeneous user modeling with generative recommendation, enabling the model to understand heterogeneous interactions and generate target items across multiple domains. 3) Enhancing the model’s ability to capture and model heterogeneous information in long sequences should also be a key focus, as longer user interaction histories often contain valuable cross-domain patterns that can further improve recommendation effectiveness.

### v2

@@ 3.2. Robustness Enhancer
Domain smoothing. Directly applying domain importance during training may cause instability, particularly for sparse domains with few samples (Wen et al., 2022). Due to the scarcity of training samples, the estimated losses for these domains can exhibit high variance, leading to noisy or unreliable values that do not accurately reflect their true importance. To alleviate this, we adopt a domain smoothing strategy following (Wen et al., 2022; Lin et al., 2024e; Zhao et al., 2023), which incorporates prior information into the current domain importance estimation. We maintain a domain importance vector and introduce a regulating factor to control the update smoothness. This leads to the following optimization objective:
@@ 4.3.1. Ablation Study (RQ2)
(2): HUM w/o user token | 0.0236 | 0.0116 | 0.0143 | 0.0068 | 0.0539 | 0.0265 | 0.0149 | 0.0068 | 0.0157 | 0.0073 | 0.0104 | 0.0051 |
(3): HUM w/o user token & prompt | 0.0217 | 0.0106 | 0.0129 | 0.0063 | 0.0563 | 0.0259 | 0.0136 | 0.0067 | 0.0146 | 0.0067 | 0.0091 | 0.0042 |
(4): HUM w/o mask | 0.0239 | 0.0113 | 0.0148 | 0.0068 | 0.0667 | 0.0309 | 0.0155 | 0.0072 | 0.0141 | 0.0063 | 0.0113 | 0.0054 |
@@
4.3.2. Compression under heterogeneity (RQ2)
To validate the effectiveness of compression enhancer in modeling heterogeneous user sequences, we design three experiments for analysis.
Domain heterogeneity experiment. We aim to evaluate the performance of HUM under domain-level heterogeneity, which refers to whether the items within a sequence originate from a single domain or multiple domains. Specifically, we divide the test data into two groups: (0) “w/o h.”, indicates that the user sequences do not contain interactions from other domains. (1) “w/ h.”, indicates that the user sequences contain interactions from other domains. We evaluate both LLM-Rec and HUM on these two groups separately. The average results across the six domains are shown in Figure 6(a). From the results, we observe the following: 1) Regardless of whether the sequences contain multi-domain interactions, HUM consistently outperforms LLM-Rec. 2) On sequences with multi-domain interactions, HUM shows a substantial improvement over LLM-Rec, which further validates HUM’s strong performance under domain-level heterogeneity and further highlight its advantage in compressing heterogeneous user sequences and capturing users’ diverse preferences.
Length heterogeneity experiment. We further evaluate the performance of HUM under sequential heterogeneity, which refers to differences in sequence length, with longer sequences generally implying greater diversity and complexity. Specifically, we divide the test data into three groups based on the number of items in each sequence: (0) “[1,3]”, indicates sequences with a length between 1 and 3. (1) “[4,6]”, indicates sequences with a length between 4 and 6; (2) “[7,10]”, indicates sequences with a length between 7 and 10. Similarly, we evaluate both LLM-Rec and HUM on these three groups separately. The average results across the six domains are shown in Figure 6(b). From the results, we observe the following: 1) HUM consistently achieves higher performance than LLM-Rec across all length groups. 2) Compared to LLM-Rec, HUM delivers more stable results in groups with longer sequences, demonstrating HUM’s effectiveness in handling length-based heterogeneity.
Compression attention experiment. We aim to investigate the impact of attention design in the compression module, which is studied by comparing unidirectional attention with bidirectional attention. Specifically, we consider the following two variants for comparison: (0) “HUM w/ bidirectional atten.”, which replaces the unidirectional attention in HUM with bidirectional attention; (1) “HUM”, our proposed method. From Table 3, we can observe that unidirectional attention achieves better performance than bidirectional one, demonstrating the advantage of our design in effectively condensing user interactions and capturing diverse preferences.
Method | R@5 | R@10 | N@5 | N@10 |
HUM w/ bidirectional atten. | 0.0119 | 0.0211 | 0.0067 | 0.0097 |
HUM | 0.0145 | 0.0247 | 0.0085 | 0.0116 |
4.3.3. Robustness in seen domains (RQ2)
@@ 4.3.4. Noise resistance(RQ2)
4.3.5. Generalization to unseen domains (RQ3)
@@ 4.3.7. Effect of mask ratio
To investigate the impact of the mask ratio on the accuracy of HUM, we vary the ratio from to and present the results in Figure 10. It is observed that 1) Moderately increasing can improve performance, as the masking mechanism helps the model leverage useful heterogeneous information from other domains while filtering noisy interactions, enhancing recommendation in the target domain. 2) However, excessively increasing may inversely hurt performance. This could be because an overly high mask ratio leads to information loss, resulting in suboptimal learning of target domain information. 3) These observations suggest that there is a sweet spot for , where the benefits of cross-domain information outweigh the costs of missing target domain details.
@@ 5. Related Work
LLM-based Recommendation. Large Language Models (LLMs) have shown exceptional success in various recommendation tasks (Wang et al., 2024; Fu et al., 2024; Ma et al., 2025; Bao et al., 2023; Deng et al., 2025; Zhao et al., 2025; Xu et al., 2025). There are two primary methods for using LLMs in recommendation: 1) LLM-based recommender, that directly employs LLMs as the recommendation models (Li et al., 2024a; Zhang et al., 2023; Lin et al., 2024a; Bao et al., 2023; Lin et al., 2024b; Prakash et al., 2023; Zheng et al., 2024; Lin et al., 2025; Li et al., 2024b). 2) LLM-enhanced recommender, while utilize LLMs for data augmentation (Wei et al., 2024; Xi et al., 2023; Liu et al., 2024) and representation (Qiu et al., 2021; Ren et al., 2024; Li et al., 2023; Tang et al., 2023). Leveraging LLMs for representation refers to generating embedding for entities(e.g., users or items) through LLMs. For instance, LLM-Rec (Tang et al., 2023) uses BERT (Devlin et al., 2018), OPT (Zhang et al., 2022) and Flan-T5 (Chung et al., 2024) to create user and item embedding based on their semantic information(i.e., plain text). In this work, we focus on leveraging LLMs for representing heterogeneous user interactions, thus we do not dive deeper into the discussion of representation methods in other fields of recommendation.
@@ 6. Conclusion and Future Work
This work emphasizes the importance of heterogeneous user modeling in LLM-based open-domain recommendation and suggests several promising directions for future research: 1) Future research may investigate theoretical foundations for compression mechanisms in heterogeneous user modeling, to better understand and justify their effectiveness from an analytical perspective. 2) Designing compression-aware architectures for large language models presents an important direction, aiming to reduce inference overhead while preserving the ability to model heterogeneous user information effectively. 3) Another direction is to explore how to combine heterogeneous user modeling with generative recommendation, enabling the model to understand heterogeneous interactions and generate target items across multiple domains.